# LCQ: Low-Rank Codebook based Quantization for Large Language Models

### v1

@@ 3.3.2 Reparameterization of Quantization Parameters
During the gradient descent process, we find it challenging to directly learn the scaling vectors $\mathcal{S}$ for all layers in a block. The main reason is that the values can vary significantly among different layers, and even within the same layer (e.g., in the fc1 layer), as shown in Figure 2. Directly learning $\mathcal{S}$ can result in excessive gradient updates for some weights while being too small for others, leading to difficulty in convergence. As in OmniQuant [28], we propose a reparameterization strategy to address the divergence issue caused by the large differences in $\mathcal{S}$ values. The reparameterization of $\mathbf{S}_{*,i}$ for the $i$th group of weights can be written as follows:
$\mathbf{S}_{*,i}$ $=g_{S}(\bar{\mathbf{S}}_{*,i},\mathbf{W}_{i})=\text{tanh}(\bar{\mathbf{S}}_{*,i})\left(\frac{\text{max}(\mathbf{W}_{i})-\text{min}(\mathbf{W}_{i})}{2}\right).$ (8)
@@
For the reparameterization of the quantization offsets $\mathbf{B}$, we employ a similar method as that for the reparameterization of $\mathbf{S}$. But we need to solve an extra issue. More specifically, according to previous studies [17, 9], it is generally necessary to include the value 00 in the set of quantized values. The main reason is that many weights in the model are close to 00, and hence excluding 00 from the quantized values would lead to significant quantization loss and severely degrade the accuracy of the final model. Assume the current offsets are $\mathbf{B}^{\prime}$. For the offsets $\mathbf{B}^{\prime}_{i}$ of the $i$th group, the reparameterized form is as follows:
@@
Although the reparameterization strategy might result in a codebook in which the maximum range could exceed the maximum range of the group weights, we can constrain the initial maximum range of the codebook to be within the maximum range of the group weights during initialization. Subsequent optimization during gradient descent will then fine-tune the codebook within the vicinity of the initial values. The training and inference process of LCQ with reparameterization is shown in Figure 3.
@@ 3.3.3 Initialization of Quantization Parameters
We use the results of AWQ [17] to initialize rank1 quantization parameters. From the results of AWQ, we can obtain the parameters $\mathbf{S}_{1}$, $\mathbf{V}_{1}$, $\mathbf{B}^{\prime}$. Specifically, AWQ [17] optimizes the equivalent transformation parameters and the clipping parameters of each linear layer. For the equivalent transformation parameters, we apply the optimized equivalent transformation parameters of AWQ to scale the weights of each linear layer. For the clipping parameters, AWQ searches for the clipping parameters for each group of weights and then clips the weights. The range and mid-point of the clipped weights correspond to $\mathbf{S}_{1}$, $\mathbf{B}^{\prime}$. In addition, $\mathbf{V}_{1}$ is a vector with uniform interval, i.e., $\mathbf{V}_{1}=[-1,-1+\frac{2}{2^{b}-1},\cdots,1-\frac{2}{2^{b}-1},1]$ in AWQ.
@@ 3.4 Double Quantization
When $N_{D}>1$, LCQ will have more parameters than rank-one codebook based methods like AWQ, which will result in extra storage cost. To mitigate the extra storage cost associated with the enlarged $\mathbf{S}$ and $\mathbf{V}$, we adopt a double quantization strategy as in QLoRA [6] to further quantize $\mathbf{S}$ and $\mathbf{V}$. Taking $\mathbf{S}$ as an example, with double quantization, the optimization goal is to minimize the quantization error of $\mathbf{S}$:
@@
The whole optimization procedure for LCQ is outlined in Algorithm 1. We sequentially quantize each Transformer block weights $\mathcal{W}_{b}$.
@@
4 Experiments
4.1 Experimental Setup
Calibration Dataset: As in AWQ [17], 128 samples randomly selected from Pile dataset [10] are used as the calibration dataset. The sequence length of each sample is 2048.
Quantization Hyperparameters: We adopt group quantization with two settings for group size: "G128" denotes a group size of 128, and "G-1" denotes that the group size is equal to the size of the channel. $N_{D}$ is set to 2 by default.
@@ 4.2 Comparison with Baselines
Table 1: Natural language generation ability on LLaMA, with perplexity as the metric. The best is shown in bold.
Config Method Retention Rate LLaMA-7B LLaMA-13B
WikiText2
fp16 - 1 5.68 5.09
W2 G128 AWQ [17] 0.134 12.90 9.73
OmniQuant [28] 0.134 12.33 9.24
LCQ 0.138 9.14 7.40
W3 G-1 AWQ [17] 0.188 8.43 6.37
OmniQuant [28] 0.188 6.73 5.91
LCQ 0.188 6.41 5.60
W3 G128 AWQ [17] 0.197 6.31 5.50
OmniQuant [28] 0.197 6.30 5.53
LCQ 0.202 6.17 5.49
C4
fp16 - 1 7.34 6.80
W2 G128 AWQ [17] 0.134 16.55 12.69
OmniQuant [28] 0.134 14.79 11.60
LCQ 0.138 11.44 9.76
W3 G-1 AWQ [17] 0.188 11.13 8.37
OmniQuant [28] 0.188 8.73 7.67
LCQ 0.188 8.35 7.42
W3 G128 AWQ [17] 0.197 8.17 7.34
OmniQuant [28] 0.197 8.18 7.35
LCQ 0.202 8.02 7.26
Table 2: Zero-shot performance of the quantized model for LLaMA-7B, with accuracy as the metric. The best is shown in bold.
Config Method PIQA WinoGrande ARC_easy
fp16 - 78.40 66.93 67.34
W2 G128 AWQ [17] 69.21 55.80 52.74
OmniQuant [28] 66.92 53.20 49.37
LCQ 72.36 60.30 59.13
W3 G128 AWQ [17] 76.33 65.19 65.07
OmniQuant [28] 76.33 64.96 64.44
LCQ 77.31 66.38 67.13
We choose AWQ [17] and OmniQuant [28] as baselines for comparison, because they are the most related works and have achieved state-of-the-art performance.
We first compare LCQ with baselines on the task of natural language generation, using OPT [38] and LLaMA [29] as pre-trained models. As in AWQ [17], we use three datasets WikiText2 [23], c4 [26] and PTB [22], with the perplexity (ppl) metric for evaluating the ability of the model to predict the next word (i.e., generative capability). The retention rate, defined as the rate between the remaining parameters to the original parameters, reflects the degree of model compression. The higher the retention rate is, the lower the degree of model compression will be. Table 1 shows the results on the three datasets, respectively. “fp16" denotes the original non-compressed model with 16-bit full-precision representation. “W2 G128” denotes a 2-bit weight quantization with a group size of 128. Other notations are defined similarly. We can find that LCQ achieves better accuracy than the baselines. In particular, in the case of 2-bit quantization, LCQ achieves significantly better accuracy than baselines. Furthermore, LCQ has a comparable retention rate with baselines, which means that the increase of $N_{D}$ brings a negligibly extra storage cost. We provide experiments on OPT models in the appendix.
We also compare LCQ with baselines on zero-shot tasks using three widely used datasets PIQA [2], WinoGrande [27] and ARC_easy [4]. Table 2 shows the accuracy of the quantized model for LLaMA-7B. We can find that LCQ achieves better accuracy than baselines on zero-shot tasks for all datasets. In particular, in the case of 2-bit quantization, the improvement is significant compared to baselines. The performance on zero-shot tasks also verifies that our LCQ method does not over-fit the calibration dataset and can maintain the generalization capability of the original model on zero-shot tasks.
5 Hyperparameter Sensitivity
First, we investigate the impact of different initialization methods for LCQ. We only perform experiments about the initialization for $\mathbf{V}_{2}$, and fix the initialization of other quantization parameters as described in Section 3.3.3. We compare three initialization methods: rand, uniform and normal. rand adopts sorted random values from $\text{Uniform}(-0.1,0.1)$ for initialization. uniform adopts the equidistant quantiles of a uniform distribution $\text{Uniform}(-1,1)$ for initialization. normal adopts the equidistant quantiles of a Gaussian distribution $\mathcal{N}(0,1)$  [6] for initialization. Figure 4(a) shows the results. We can find that different initialization methods do not have a significant impact on the performance. Hence, LCQ is robust to the initialization methods for $\mathbf{V}_{2}$. By default, we use normal for initialization to avoid extra hyperparameter settings.
@@ 5 Hyperparameter Sensitivity
Figure 4: Hyperparameter sensitivity with “W2 G128" for OPT-1.3B.
The result about $N_{G}$ is in Figure 5(a). The smaller the $N_{G}$, the finer the quantization granularity. In particular, $N_{G}=\frac{N_{W}}{G}$ corresponds to the coarsest granularity, indicating that the whole linear layer corresponds a single $\mathbf{V}$. We can find that reducing $N_{G}$ can bring a slight improvement in model accuracy but also introduces some additional storage cost. To balance model accuracy and storage cost, we set $N_{G}=32$ by default.
The number of bits for double quantization will also affect the performance. Figure 5(b) shows the result. In particular, $\mathbf{S}_{1}$ is not quantized and kept in fp16, consistent with baseline methods. We can find that when using 4-bit quantization for $\mathbf{S}$ and 8-bit quantization for $\mathbf{V}$, the accuracy loss is minor or even negligible compared to methods without double quantization. However, when the number of quantization bits for $\mathbf{V}$ is reduced to 4, the accuracy loss is more significant. Hence, we use 4-bit quantization for $\mathbf{S}$ and 8-bit quantization for $\mathbf{V}$ by default. Double quantization is performed in groups of 16 elements by default.
(a) $N_{G}$.
(b) Double quantization.
Figure 5: Hyperparameter sensitivity with “W2 G128" for OPT-1.3B.
We also explore the impact of $N_{D}$ on model performance, the result of which is shown in Table 3. We can find that compared to the results with $N_{D}=1$, the results with $N_{D}=2$ improve significantly, which verifies that increasing the rank of the codebook can lead to better accuracy. Further increasing $N_{D}$ can further improve the accuracy, but will also incur more storage cost. Hence, we set $N_{D}$ to 2 by default. Furthermore, we can also find that although the optimized parameters of AWQ are used to initialize $\mathbf{S}_{1},\mathbf{V}_{1},\mathbf{B}$ in LCQ, further learning (optimizing) these parameters can achieve better accuracy than fixing them. We provide more hyperparameter sensitivity analysis in the appendix.
Table 3: Perplexity on WikiText2 when using different $N_{D}$ and whether fixing $\mathbf{S}_{1},\mathbf{B},\mathbf{V}_{1}$, with “W2 G128" for OPT-1.3B.
@@ 6 Conclusion
In this paper, we propose a novel weight quantization method called LCQ for compressing LLMs. Experiments show that LCQ can outperform existing methods to achieve the best performance.
Limitations. We have only conducted experiments on uni-modal models and have not yet validated our method on multi-modal models. We will validate our method on multi-modal models in the future.
Appendix A Training Details
We perform experiments with PyTorch [25] on an RTX A6000 GPU card of 48GB memory, and use mixed-precision optimization to reduce the memory and computational cost during training. The ADAMW optimizer [21] is used, with an initial learning rate set to 0.01, and the weight decay factor set to 0. The learning rate is gradually reduced using a cosine annealing schedule [20] over a total of 10 training epochs. The batch size for training is set to 4, and gradient accumulation is employed to prevent memory overflow. The linear layer weights of all Transformer blocks are quantzied. The embedding layer and the output layer are not quantized.
Appendix B Experiments on OPT Models
Table 4 shows the experiments on OPT models.
Table 4: Natural language generation ability on OPT, with perplexity as the metric.
Config Method Retention Rate OPT-125M OPT-1.3B OPT-6.7B
WikiText2
fp16 - 1 27.65 14.62 10.86
W2 G128 AWQ [17] 0.134 340.56 42.72 16.73
OmniQuant [28] 0.134 117.25 34.69 17.15
LCQ 0.138 68.75 25.28 15.49
W3 G-1 AWQ [17] 0.188 57.46 24.67 16.54
OmniQuant [28] 0.188 41.50 18.85 12.98
LCQ 0.188 36.88 17.54 11.96
W3 G128 AWQ [17] 0.197 35.99 16.41 11.36
OmniQuant [28] 0.197 35.23 16.37 11.53
LCQ 0.202 32.47 15.97 11.36
C4
fp16 - 1 26.56 16.07 12.71
W2 G128 AWQ [17] 0.134 246.47 40.63 19.55
OmniQuant [28] 0.134 82.47 31.59 18.34
LCQ 0.138 54.20 25.31 16.93
W3 G-1 AWQ [17] 0.188 48.83 24.55 17.79
OmniQuant [28] 0.188 36.36 19.82 15.20
LCQ 0.188 33.81 18.76 13.81
W3 G128 AWQ [17] 0.197 32.63 17.80 13.41
OmniQuant [28] 0.197 31.92 17.69 13.43
LCQ 0.202 30.12 17.30 13.30
PTB
fp16 - 1 38.99 20.29 15.77
W2 G128 AWQ [17] 0.134 516.64 73.33 25.95
OmniQuant [28] 0.134 163.66 49.46 24.52
LCQ 0.138 89.73 38.12 22.26
W3 G-1 AWQ [17] 0.188 86.42 34.65 21.56
OmniQuant [28] 0.188 58.22 26.77 20.48
LCQ 0.188 53.72 25.38 17.26
W3 G128 AWQ [17] 0.197 50.38 23.23 16.78
OmniQuant [28] 0.197 48.78 22.85 16.76
LCQ 0.202 46.75 22.40 16.65

### v2

@@ 3.3.2 Reparameterization of Quantization Parameters
During the gradient descent process, we find it challenging to directly learn the scaling vectors $\mathcal{S}$ for all layers in a block. The main reason is that the values can vary significantly among different layers, and even within the same layer (e.g., in the fc1 layer), as shown in Figure 2. Directly learning $\mathcal{S}$ can result in excessive gradient updates for some weights while being too small for others, leading to difficulty in convergence. As in OmniQuant (Shao et al., 2024), we propose a reparameterization strategy to address the divergence issue caused by the large differences in $\mathcal{S}$ values. The reparameterization of $\mathbf{S}_{*,i}$ for the $i$th group of weights can be written as follows:
$\begin{split}\mathbf{S}_{*,i}&=g_{S}(\bar{\mathbf{S}}_{*,i},\mathbf{W}_{i})\\ &=\text{tanh}(\bar{\mathbf{S}}_{*,i})\left(\frac{\text{max}(\mathbf{W}_{i})-\text{min}(\mathbf{W}_{i})}{2}\right).\end{split}$ (8)
@@
For the reparameterization of the quantization offsets $\mathbf{B}$, we employ a similar method as that for the reparameterization of $\mathbf{S}$. But we need to solve an extra issue. More specifically, according to previous studies (Lin et al., 2024; Frantar et al., 2023), it is generally necessary to include the value 00 in the set of quantized values. The main reason is that many weights in the model are close to 00, and hence excluding 00 from the quantized values would lead to significant quantization loss and severely degrade the accuracy of the final model. Assume the current offsets are $\mathbf{B}^{\prime}$. For the offsets $\mathbf{B}^{\prime}_{i}$ of the $i$th group, the reparameterized form is as follows:
@@
Although the reparameterization strategy might result in a codebook in which the maximum range could exceed the maximum range of the group weights, we can constrain the initial maximum range of the codebook to be within the maximum range of the group weights during initialization.
Subsequent optimization during gradient descent will then fine-tune the codebook within the vicinity of the initial values. The training and inference process of LCQ with reparameterization is shown in Figure 3.
@@ 3.3.3 Initialization of Quantization Parameters
We use the results of AWQ (Lin et al., 2024) to initialize rank-1 quantization parameters. From the results of AWQ, we can obtain the parameters $\mathbf{S}_{1}$, $\mathbf{V}_{1}$, $\mathbf{B}^{\prime}$. Specifically, AWQ (Lin et al., 2024) optimizes the equivalent transformation parameters and the clipping parameters of each linear layer. For the equivalent transformation parameters, we apply the optimized equivalent transformation parameters of AWQ to scale the weights of each linear layer. For the clipping parameters, AWQ searches for the clipping parameters for each group of weights and then clips the weights. The range and mid-point of the clipped weights correspond to $\mathbf{S}_{1}$, $\mathbf{B}^{\prime}$. In addition, $\mathbf{V}_{1}$ is a vector with uniform interval, i.e., $\mathbf{V}_{1}=[-1,-1+\frac{2}{2^{b}-1},\cdots,1-\frac{2}{2^{b}-1},1]$ in AWQ.
@@ 3.4 Double Quantization
When $N_{D}>1$, LCQ will have more parameters than rank-one codebook based methods like AWQ, which will result in extra storage cost. To mitigate the extra storage cost associated with the enlarged $\mathbf{S}$ and $\mathbf{V}$, we adopt a double quantization strategy as in QLoRA (Dettmers et al., 2023) to further quantize $\mathbf{S}$ and $\mathbf{V}$. Taking $\mathbf{S}$ as an example, with double quantization, the optimization goal is to minimize the quantization error of $\mathbf{S}$:
@@
The whole optimization procedure for LCQ is outlined in Algorithm 1. We sequentially quantize the weights $\mathcal{W}_{b}$ of each Transformer block.
4 Experiments
4.1 Experimental Setup
@@
As in AWQ (Lin et al., 2024), 128 samples randomly selected from Pile dataset (Gao et al., 2021) are used as the calibration dataset. The sequence length of each sample is 2048.
We adopt group quantization with two settings for group size: “G128” denotes a group size of 128, and “G-1” denotes that the group size is equal to the size of the channel. $N_{D}$ is set to 2 by default.
Table 1: Natural language generation ability on OPT, LLaMA and LLaVA models, with perplexity on WikiText2 as the metric. The best is shown in bold. The retention rate reflects the degree of model compression.
fp16 W2 G128 W3 G-1 W3 G128
AWQ Omniquant LCQ AWQ Omniquant LCQ AWQ Omniquant LCQ
Retention Rate 1 0.134 0.134 0.138 0.188 0.188 0.188 0.197 0.197 0.202
OPT-125M 27.65 340.56 117.25 68.75 57.46 41.50 36.88 35.99 35.23 32.47
OPT-1.3B 14.62 42.72 34.69 25.28 24.67 18.85 17.54 16.41 16.37 15.97
OPT-6.7B 10.86 16.73 17.15 15.49 16.54 12.98 11.96 11.36 11.53 11.36
LLaMA-7B 5.68 12.90 12.33 9.14 8.43 6.73 6.41 6.31 6.30 6.17
LLaMA-13B 5.09 9.73 9.24 7.40 6.37 5.91 5.60 5.50 5.53 5.49
LLaVA-7B 6.83 23.63 36.58 12.29 20.97 8.74 8.05 7.87 7.86 7.50
LLaVA-13B 5.98 11.89 18.08 9.27 8.28 6.99 6.65 6.57 6.66 6.52
Table 2: Natural language generation ability on OPT, LLaMA and LLaVA models, with perplexity on c4 as the metric. The best is shown in bold. The retention rate reflects the degree of model compression.
fp16 W2 G128 W3 G-1 W3 G128
AWQ Omniquant LCQ AWQ Omniquant LCQ AWQ Omniquant LCQ
Retention Rate 1 0.134 0.134 0.138 0.188 0.188 0.188 0.197 0.197 0.202
OPT-125M 26.56 246.47 82.47 54.20 48.83 36.36 33.81 32.63 31.92 30.12
OPT-1.3B 16.07 40.63 31.59 25.31 24.55 19.82 18.76 17.80 17.69 17.30
OPT-6.7B 12.71 19.55 18.34 16.93 17.79 15.20 13.81 13.41 13.43 13.30
LLaMA-7B 7.34 16.55 14.79 11.44 11.13 8.73 8.35 8.17 8.18 8.02
LLaMA-13B 6.80 12.69 11.60 9.76 8.37 7.67 7.42 7.34 7.35 7.26
LLaVA-7B 9.27 27.83 33.54 14.74 24.66 12.11 11.04 10.70 10.79 10.32
LLaVA-13B 8.26 16.42 20.35 12.02 11.46 9.84 9.37 9.12 9.26 9.03
We perform experiments with PyTorch (Paszke et al., 2019) on an RTX A6000 GPU card of 48GB memory, and use mixed-precision optimization to reduce the memory and computational cost during training. The ADAMW optimizer (Loshchilov & Hutter, 2019) is used, with an initial learning rate set to 0.01, and the weight decay factor set to 0. The learning rate is gradually reduced using a cosine annealing schedule (Loshchilov & Hutter, 2017) over a total of 10 training epochs. The batch size for training is set to 4, and gradient accumulation is employed to prevent memory overflow. The weights of linear layer in all Transformer blocks are quantzied. The embedding layer and the output layer are not quantized.
@@ 4.2 Comparison with Baselines
We choose AWQ (Lin et al., 2024) and OmniQuant (Shao et al., 2024) as baselines for comparison, because they are the most related works and have achieved state-of-the-art performance.
We first compare LCQ with baselines on the task of natural language generation, using OPT (Zhang et al., 2022), LLaMA (Touvron et al., 2023) and LLaVA (Liu et al., 2022a) as pre-trained models. As in AWQ (Lin et al., 2024), we use datasets WikiText2 (Merity et al., 2017) and c4 (Raffel et al., 2020), with the perplexity (ppl) metric for evaluating the ability of the model to predict the next word (i.e., generative capability). The retention rate, defined as the rate between the remaining parameters to the original parameters, reflects the degree of model compression. The higher the retention rate is, the lower the degree of model compression will be. Table 1 shows the perplexity on WikiText2 dataset, and Table 2 shows the perplexity on c4 dataset. “fp16” denotes the original non-compressed model with 16-bit full-precision representation. “W2 G128” denotes a 2-bit weight quantization with a group size of 128. Other notations are defined similarly. We can find that LCQ achieves better accuracy than the baselines. In particular, in the case of 2-bit quantization, LCQ achieves significantly better accuracy than baselines. This significant improvement in accuracy demonstrates the effectiveness of the low-rank codebook in capturing the nuances of language modeling tasks. Furthermore, LCQ has a comparable retention rate with baselines, which means that the increase of $N_{D}$ brings a negligibly extra storage cost.
@@
(c) $N_{G}$.
(d) Double quantization.
Figure 4: Hyperparameter sensitivity with “W2 G128” for OPT-1.3B.
Table 3: Zero-shot performance of the quantized model for LLaMA-7B, with accuracy as the metric. The best is shown in bold. ARC denotes ARC_easy.
Config Method PIQA WinoGrande ARC
fp16 - 78.40 66.93 67.34
W2 G128 AWQ 69.21 55.80 52.74
OmniQuant 66.92 53.20 49.37
LCQ 72.36 60.30 59.13
W3 G128 AWQ 76.33 65.19 65.07
OmniQuant 76.33 64.96 64.44
LCQ 77.31 66.38 67.13
In addition to perplexity, we also compare LCQ with baselines on zero-shot tasks using three widely used datasets PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2020) and ARC_easy (Clark et al., 2018). Table 3 shows the accuracy of the quantized model for LLaMA-7B. We can find that LCQ achieves better accuracy than baselines on zero-shot tasks for all datasets. In particular, in the case of 2-bit quantization, the improvement is significant compared to baselines. The performance on zero-shot tasks also verifies that our LCQ method does not over-fit the calibration dataset and can maintain the generalization capability of the original model on zero-shot tasks.
4.3 Hyperparameter Sensitivity
First, we investigate the impact of different initialization methods for LCQ. We only perform experiments about the initialization for $\mathbf{V}_{2}$, and fix the initialization of other quantization parameters as described in Section 3.3.3. We compare three initialization methods: rand, uniform and normal. rand adopts sorted random values from $\text{Uniform}(-0.1,0.1)$ for initialization. uniform adopts the equidistant quantiles of a uniform distribution $\text{Uniform}(-1,1)$ for initialization. normal adopts the equidistant quantiles of a Gaussian distribution $\mathcal{N}(0,1)$  (Dettmers et al., 2023) for initialization. Figure 4(a) shows the results. We can find that different initialization methods do not have a significant impact on the performance. Hence, LCQ is robust to the initialization methods for $\mathbf{V}_{2}$. By default, we use normal for initialization to avoid extra hyperparameter settings.
Table 4: Perplexity on WikiText2 when using different $N_{D}$ and whether fixing $\mathbf{S}_{1},\mathbf{B},\mathbf{V}_{1}$, with “W2 G128” for OPT-1.3B.
@@ 4.3 Hyperparameter Sensitivity
The result about $N_{G}$ is in Figure 4(c). The smaller the $N_{G}$ is, the finer the quantization granularity will be. In particular, $N_{G}=\frac{N_{W}}{G}$ corresponds to the coarsest granularity, indicating that the whole linear layer corresponds a single $\mathbf{V}$. We can find that reducing $N_{G}$ can bring a slight improvement in model accuracy but also introduces some additional storage cost. To balance model accuracy and storage cost, we set $N_{G}=32$ by default.
The number of bits for double quantization will also affect the performance. Figure 4(d) shows the result. In particular, $\mathbf{S}_{1}$ is not quantized and kept in fp16, consistent with baseline methods. We can find that when using 4-bit quantization for $\mathbf{S}$ and 8-bit quantization for $\mathbf{V}$, the accuracy loss is minor or even negligible compared to methods without double quantization. However, when the number of quantization bits for $\mathbf{V}$ is reduced to 4, the accuracy loss is more significant. Hence, we use 4-bit quantization for $\mathbf{S}$ and 8-bit quantization for $\mathbf{V}$ by default. Double quantization is performed in groups of 16 elements by default.
We also explore the impact of $N_{D}$ on model performance, the result of which is shown in Table 4. We can find that compared to the results with $N_{D}=1$, the results with $N_{D}=2$ improve significantly, which verifies that increasing the rank of the codebook can lead to better accuracy. Further increasing $N_{D}$ can further improve the accuracy, but will also incur more storage cost. Hence, we set $N_{D}$ to 2 by default. Furthermore, we can also find that although the optimized parameters of AWQ are used to initialize $\mathbf{S}_{1},\mathbf{V}_{1},\mathbf{B}$ in LCQ, further learning (optimizing) these parameters can achieve better accuracy than fixing them.
@@ 5 Conclusion
In this paper, we propose a novel weight quantization method, called low-rank codebook based quantization (LCQ), for large language models (LLMs). LCQ leverages a low-rank codebook with a rank potentially greater than one, significantly enhancing the accuracy of quantized models while incurring only a negligible additional storage cost. Experiments demonstrate that LCQ surpasses existing methods, especially in low bit-width scenarios. It effectively reduces storage and computational requirements while maintaining high accuracy, making LCQ a practical solution for deploying large models on resource-constrained devices.
Impact Statement
The LCQ method represents a significant advancement in the field of model compression, offering a practical solution to the challenges of deploying large language models in resource-constrained environments. By enabling high compression ratios with minimal accuracy loss, LCQ makes it feasible to deploy LLMs on mobile and edge devices, opening up new possibilities for real-time applications in natural language processing and multimodal tasks. Furthermore, the reduced energy consumption associated with LCQ contributes to the development of sustainable AI technologies, addressing the growing environmental concerns related to large-scale AI model deployment. The compatibility of LCQ with existing quantization techniques ensures its versatility and broad applicability, making it a valuable tool for both academia and industry.